OpenStudy (anonymous):

a baseball pitcher throws a baseball with a speed of 41m/s. Estimate the average acceleration of the ball during the throwing motion. In throwing the baseball, the pitcher accelerates the ball through a displacement of 3.5m... I worked it out but my answer is wrong and I don't know why..

OpenStudy (anonymous):

it is 240.14m/s^2 isn't it?

OpenStudy (anonymous):

yes but how? you are like a physics genius..can i borrow your brain for my exam?

OpenStudy (anonymous):

lol, you make me blush! You see, the pitcher has accelerated the ball from rest to 41m/s in a distance of 3.5m So, using

OpenStudy (anonymous):

v=u+at -(1) s=(v+u)1/2t -(2) So ans: 240.1428ms^-2

OpenStudy (anonymous):

\[v _{f}^{2}=v _{0}^{2}+2as\] calculate 'a'

OpenStudy (anonymous):

@Mif.100 how did you find the time?

OpenStudy (anonymous):

@rajathsbhat and then after that i can calculate time with v=vo x at?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

@rajathsbhat didn't you leave out the displacement part? v^2=vi^2+2a(x-xi)

OpenStudy (anonymous):

in v=u+at

OpenStudy (anonymous):

s=displacement

OpenStudy (anonymous):

oo okay, thank you! :) i have two more questions.. will post them soon! x

OpenStudy (anonymous):

welcm

OpenStudy (anonymous):

looking forward to it :)